Mary expects XYZ stock price will differ a lot from $80 but not sure whether it will be smaller or larger. She wants to create a straddle. Calculate her initial cash flow.

Strike price Call option price Put option price

$80 . $8.16 . $3.39

$90 $6.98 $5.35

If it's a cash inflow, enter a positive number. If it's a cash outflow, enter a negative number.
Answer #1

Solution:

Straddle strategy is formed when a call and a put option is purchase for the same strike price. When movement of the stock is high ( Either on lower side or higher side then this strategy will give profit ).

Straddle 1:

Buying a put and a call option of $80 strike .

Total initial cashflow = - $8.16 -$3.39 = -$11.55

This strategy will generate profit when share price falls below $80 - $11.55 = 68.45 or share price goes beyond $80 +11.55 = 91.55

Straddle 2 :

Buying a put and a call options of $90 strike price

Total initial cash flow = -6.98 -5.35 = -$12.33

This strategy will generate profit when share falls below $90 - 12.33 = $77.67 or it goes above $90 + $12.33 = $102.33

So answer is -$11.55 and -$12.33 in straddle strategy for $80 and $90 strike price respectively